OpenStudy (anonymous):

Evaluate: -5.3 + (-4.5) =

OpenStudy (anonymous):

-9.8

OpenStudy (anonymous):

-5.3+(-4.5) = -5.3-4.5 = -9.8

OpenStudy (anonymous):

negative plus negative is always negative, so bascically all you need to do is add them up and add a negative sign

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

no probelem